Sol Summary – November 1st

Sol: 3
Summary Title: Resting Day
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 9 a.m., after a deserved rest. On Sol 2, the crew worked hard and ended the day with a small “Halloween party” with pizza, mocktails, and Latin music.
The day was focused on the projects, preparing future EVAs
Look Ahead Plan: Continue with the project planning of the EVAs and be prepared for tomorrow´s media events.
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: Nominal
EVA: No EVA today
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report.
Support Requested: We need to add a new email address to the mailing list of mission support because our HSO is having issues sending emails from her current email to mission support; the new email we need to add is andrea10stars1g. For the media stream with Badalona Comunicacio, I informed the Media and Public Relations Director and MDRS Director we need an internet connection tomorrow between 8 a.m. and 2 p.m.

Operations Report – November 1st

SOL: 3
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: No
Currently charging: Yes
Opportunity rover used: No
Currently charging: Yes
Curiosity rover used: No
Currently charging: Yes
Perseverance rover used: No
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: 23 gallons
Water (static tank): 406 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): off
Toilet tank emptied: yes, the indicator only got to “⅓” instead of “empty”
Summary of internet: No comments
Summary of suits and radios: No comments
Summary of GreenHab operations: fertilization, thermostat adjusted to 65 F
WATER USE: 3.6 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: used the blender, 3 large and 1 small paper filters, the microwave, the oven, 2 large erlenmeyer flasks, 1 large and 4 small beaker glasses, the scale, and the mortar and pestle. All this was used to recover starch from rice and potatoes and then mix it with a soil sample to create a concrete-like material.
Dual split: Off
Summary of RAM operations: Not used today
Summary of any observatory issues: “Mount error” was fixed by the observatory director, however 2 hs later the “Camera error” showed up again, which stopped the image acquisition.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: None

Journalist Report – November 1st

by Marina Buqueras Rovira


¡El que persevera, alcanza!

Sol 3

La noche terminó con los sentimientos al máximo, el agotamiento nos arropaba aunque la satisfacción de lo conseguido nos hacía vibrar de felicidad, ¡las risas nos hacen hasta llorar!

¡Amaneció tranquilo…, el amanecer aquí es tan bonito!

9:30 am Desayunamos y empezamos a poner en marcha cada uno de los trabajos planificados para hoy.

La limpieza general en el Hab entre todos, Andrés acabando el inventario de los alimentos (nos tiene las galguerías, golosinas, dulces, chucherías… bien restringidas), Luis concentrado y obteniendo imágenes inimaginables del cielo diariamente, Andrea se prepara para iniciar las mezclas y obtener el almidón en el ScienceDome, Tomás con los receptores gps…

La mañana pasa muy rápido, tan concentrados…

14:30 pm Tom y Marina cocinaron a dúo el almuerzo para la tripulación, hoy tenemos brócoli con boniato y zanahoria acompañado de puré de patatas. Y el momento sobremesa nos recompone el alma.

17 pm La tarde siguió con calma pero sin pausa, Andrés se dirige al Greenhab a fertilizar y los demás enfocados en sus tareas y proyectos.

Mañana más y mejor!


The one who perseveres, achieves!

Sol 3

The night ended with emotions running high, exhaustion enveloping us even as the satisfaction of our achievements made us thrilled with happiness – laughter making us even cry!

The morning dawned peacefully; the sunrise here is so beautiful!
9:30 am We had breakfast and began to set each of today’s tasks in motion.

General cleaning in the Hab, Andrés finishing the food inventory (he has our snacks, treats, sweets, and candies well rationed), Luis focused on capturing unbelievable images of the sky daily, Andrea is getting ready to start the mixtures and obtain starch in the ScienceDome, Tomás with the GPS receivers…

The morning goes by so fast, so focused…

2:30 pm Tom and Marina cooked lunch together for the crew. Today we had broccoli with sweet potato and carrot accompanied by mashed potatoes. And the post-meal moment reassembles our souls.

5 pm The afternoon continued calmly but without pause, Andrés headed to the Greenhab to fertilize, and the others focused on their tasks.

More and better tomorrow!

GreenHab Report – November 1st

GreenHab Officer: Andres David Reina
Environmental control: (heater, fan, door open?)
Open door: 9:30 am:
-     Temperature: 23,4C
-     Humidity:10%
Close door: 18:05
-     Temperature: 22,3 C
-     Humidity: 15%
Hours of supplemental light: None
Daily water usage for crops: 1.56 gallons.
Daily water usage for research and/or other purposes : 2.1 gallons/ Science Dome operations.
Water in Blue Tank 192,94 gallons
Time(s) of watering for crops: n/a
Changes to crops: A slight growth in the tomato crop. Two existing germinations with an average size of 2 cm in height. Growth is observed in cucumber and lettuce crops.

– The Greenhab door opened at 9:30 am. Environmental measurements of temperature and relative humidity were taken. It begins with the fertilizer process for the crops. Miracle-Gro Water soluble Supplement is used, 1 tablespoon of substrate per 1.56 gallons. The crops are irrigated. The thermostat level has been set to 65F. Attached are some photos of the current state of the greenhouse.

Support/supplies needed:
– There are two crops in the growing process without identification. Please could you provide instructions on the type of fruit planted to begin with its corresponding labeling.

Astronomy Report – November 1st



Name: Luis Diaz
Crew: 285
Date: 11/01/2023
Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: NGC 281 (Pacman nebula) and maybe others.
Images submitted with this report: Orion Nebula (M41)
Problems Encountered: Just the same problems with the "Mount error" and "Camera error"; however, Peter is helping me to fix it.
IMPORTANT!!!!!!!! : Out of the entire sequence of photographs of the 5 bands, I found an anomaly in one of the photographs of the green band where 3 bright points are visible in a 3-MINUTE exposure photograph. Due to the exposure time, it doesn’t seem to be a meteor or satellites. I hope that Mission Support can assist me in delving into this and finding a possible explanation. If further information is needed on this matter, please let me know. Attached is a comparison between a photograph similar to all the others where the 3 points are not visible and the particular photo I mentioned earlier.
Solar Features Observed: –
Images submitted with this report: –
Problems Encountered: I couldn’t use the observatory due the clouds.

Copyright © The Mars Society. All rights reserved. | Main Site